find the total length of the curve r=4(1-sin theta) form theta = 90 degree to theta= 270 degree


1
Expert's answer
2022-06-02T05:04:59-0400


We can find the length of the curve by the next formula:

"l=\int_{ \theta_1}^{ \theta_2} \sqrt{(r(\theta))^2+r '(\theta))^2}d \theta."

"r(\theta)=4(1-sin\theta),"

"r'(\theta)=-4cos\theta,"

so we have:

"l=\int_{ 90\degree}^{ 270\degree} \sqrt{(4(1-sin\theta))^2+(-4cos\theta)^2}d \theta="


"=\int_{ 90\degree}^{ 270\degree} \sqrt{16-32sin\theta+16sin^2\theta+16cos^2\theta}d \theta="


"=\int_{ 90\degree}^{ 270\degree} \sqrt{16-32sin\theta+16}d \theta="


"=\int_{ 90\degree}^{ 270\degree} \sqrt{32(1-sin\theta)}d \theta="


"=4\sqrt{2}\int_{ 90\degree}^{ 270\degree} \sqrt{1-sin\theta}d \theta="


"=4\sqrt{2}\int_{ 90\degree}^{ 270\degree} \sqrt{(1-sin\theta) \frac{1+sin\theta}{1+sin\theta}}d \theta="


"=4\sqrt{2}\int_{ 90\degree}^{ 270\degree} \sqrt{ \frac{1-sin^2\theta}{1+sin\theta}}d \theta="


"=4\sqrt{2}\int_{ 90\degree}^{ 270\degree} \sqrt{ \frac{cos^2\theta}{1+sin\theta}}d \theta="


"=4\sqrt{2}\int_{ 90\degree}^{ 270\degree} \frac{-cos\theta}{\sqrt{{1+sin\theta}}}d \theta="


"=-4\sqrt{2}\int_{ 90\degree}^{ 270\degree} \frac{cos\theta}{\sqrt{{1+sin\theta}}}d \theta."


Substitution: "1+sin\theta=u," "du=cos\theta d\theta,"

"1+sin(90\degree)=1+1=2,"

"1+sin(270\degree)=1+(-1)=0."


We have:

"l=-4\sqrt{2}\int_{ 2}^{0} \frac{du}{\sqrt{u}}=4\sqrt{2} \cdot 2\sqrt{u}|_0^2=8\sqrt{2}(\sqrt{2}-0)=16."


Answer: 16.